# Enhancing Character-Level Understanding in LLMs through Token Internal Structure Learning

### v2

@@ Abstract
footnotetext: * These authors contributed equally to this work.footnotetext: † Corresponding author.
@@ 1 Introduction
For instance, LLMs frequently struggle with simple character-counting tasks. When prompted with questions like "How many ’r’s are in ’strawberry’?" (Xu and Ma, 2024), many models fail to provide the correct answer due to their limited understanding of character positions within tokens. This limitation is more pronounced in languages like Chinese, where meaning relies heavily on character composition and sequence. Models such as GPT-4o (Hurst et al., 2024) often misidentify specific character positions in tokenized text. For example, when asked to locate the character {CJK}UTF8gkai"阁" in the sentence {CJK}UTF8gkai"为什么总称呼对方为阁下?" (Why do you always address each other as ’Your Excellency’?) (Wu et al., 2023), they frequently provide incorrect positions.
@@
Figure 1: Comparison of CSC performance before and after applying TIPA, demonstrating improved character-level understanding.
@@
For example, a token like {CJK}UTF8gkai"小说" (novel) would be decomposed in TIPA as a JSON structure: {CJK}UTF8gkai{2: "说", 1: "小"}, mapping each character to its position in descending order. This approach helps the model develop a structural understanding beyond typical left-to-right reading, which is crucial for tasks requiring precise character positioning.
@@
Table 1: Impact of accurate character position prediction in CSC tasks.
@@
    Introduction of TIPA and MTIPA: Presenting TIPA and its extension, Multi-Token Internal Position Awareness (MTIPA), which improve models’ ability to capture character positions for accurate predictions.
@@ 2 Related Work
Tokenization methods like BPE (Sennrich, 2015; Wang et al., 2020) and WordPiece (Schuster and Nakajima, 2012) improve computational efficiency in LLMs but obscure internal character structures. Kaushal and Mahowald (2022) found that while larger models encode character-level details better, they may not explicitly understand character positions within tokens.
@@ 4 Redefining the Chinese Spelling Correction Task
We redefine the Chinese Spelling Correction (CSC) task to require the model to output the positions of incorrect characters along with their corrections. For example, the sentence {CJK}UTF8gkai"业内人事称撤向东南亚亦属正常" (translation: "Industry insiders say that withdrawing to Southeast Asia is also normal") becomes:
@@
UTF8gkai ["position": 4, "incorrect": "事", "correction": "士"]
Here, the character at position 4, "{CJK}UTF8gkai事" (meaning "affair"), is incorrect and should be corrected to "{CJK}UTF8gkai士" (meaning "scholar").
@@ 5 Experiments
We conducted comprehensive experiments to validate the effectiveness of TIPA and MTIPA. Our experiments are divided into two main parts corresponding to the tasks described earlier.
@@ 5.2 Baseline Models
We compared our methods against several baseline models, including Pure-SFT, GPT-4o (Hurst et al., 2024), DeepSeek v2.5, ERNIE-4.0, and GLM-4-Plus, to assess the relative performance improvements.
@@ 5.3 Training Details
We fine-tuned the open-source Qwen2.5-7B (Yang et al., 2024) model using LoRA (Hu et al., 2021) to incorporate TIPA and MTIPA. LoRA allows efficient adaptation of large models without modifying the original model weights, making it suitable for our experiments.
@@ 5.4.1 Results
Table 4: Results on position-based CSC using CSCD-NS test dataset, showing that TIPA-7B and MTIPA-7B outperform the baseline in all evaluated metrics.
@@ 5.5.1 Results
Table 5: Experiment 2: Traditional CSC results on CSCD-NS test dataset. TIPA improves character-level detection and correction, with reverse-order TIPA showing greater gains at larger model scales.
@@ 5.5.2 Evaluation on LEMON Dataset
Table 6: Experiment 2: Character-level F1 scores on LEMON and CSCD-NS for various domains. TIPA-7B consistently achieves higher F1, indicating better generalization.
@@ 6.1 Position Prediction Accuracy
Figure 3: Comparison of position prediction accuracy by character position. MTIPA-7B achieves consistently higher accuracy, especially at longer positions.
@@ 6.2 Training Dynamics
Figure 4: Training dynamics showing character-level metrics and position accuracy. MTIPA-7B achieves the highest performance ceiling.
@@ 7 Conclusion
We introduced Token Internal Position Awareness (TIPA) and Multi-Token Internal Position Awareness (MTIPA) to enhance large language models’ (LLMs) ability to accurately predict character positions within tokens. By training LLMs on reverse character prediction tasks using the tokenizer’s vocabulary, TIPA and MTIPA effectively address the limitations imposed by tokenization methods like Byte-Pair Encoding (BPE). Our experiments on Chinese Spelling Correction (CSC) tasks demonstrate that models equipped with TIPA and MTIPA achieve higher position prediction accuracy and improved overall performance. Additionally, TIPA benefits downstream tasks requiring character-level information even without explicit position prediction. This approach enriches LLMs’ internal structure awareness with minimal training overhead, offering a practical solution for tasks demanding fine-grained linguistic details.
@@ 8 Limitations
  • The experiments were conducted exclusively on the Chinese Spelling Correction task. The applicability of TIPA and MTIPA to other languages, such as Japanese or Korean, remains to be explored.
@@
  • The improvement of model capability by introducing TIPA and MTIPA was carried out under LoRA fine-tuning. It remains to be explored whether introducing TIPA and MTIPA in full training or general models can enhance model capability.
@@ A.3 Datasets
We utilized several datasets for training and evaluation. The datasets and their characteristics are summarized in Tables 12 and 13. The total number of unique tokens after deduplication across all datasets is 24,994.
@@ A.7 Loss Analysis
Figure 6: Experiment 2: Training and validation loss comparison across different methods in the traditional CSC task. The TIPA-7B model exhibits the fastest reduction in loss during training, indicating more efficient learning and better generalization capabilities compared to the baseline.

### v5

@@ Abstract
footnotetext: † Corresponding author.
@@ 1 Introduction
For instance, LLMs frequently struggle with simple character-counting tasks. When prompted with questions like "How many r’s are in strawberry’?" (Xu and Ma, 2024), many models fail to provide the correct answer due to their limited understanding of character positions within tokens, this limitation is more pronounced in languages like Chinese, where meaning relies heavily on character composition and sequence. Models such as GPT-4o (Hurst et al., 2024) often misidentify specific character positions in the tokenized text. For example, when asked to locate the character {CJK}UTF8gkai"" in the sentence {CJK}UTF8gkai"为什么称呼对方" (Why do you always address each other as ’Your Excellency’?) (Wu et al., 2023), they frequently provide incorrect positions.
@@
Figure 1: (a) demonstrates LLMs’ inability to perform spelling correction correctly without learning token-internal character order. (b) shows TIPA’s core dataset construction (left) and its character-level task enhancement through token structure understanding without architectural changes (right).
@@
For example, a token like {CJK}UTF8gkai"小说" (novel) would be decomposed in TIPA as a JSON structure: {CJK}UTF8gkai{2: "", 1: ""}, mapping each character to its position in descending order. This approach helps the model develop a structural understanding beyond typical left-to-right reading, which is crucial for tasks requiring precise character positioning.
@@
Table 1: Accurately predicting the position of erroneous characters and providing both the incorrect and corrected characters serves two purposes: the incorrect character verifies the model’s ability to precisely locate errors, while the corrected character fulfills the error correction task. This approach also reduces the number of output tokens required by the model.
@@
    Introduction of TIPA and MTIPA: Presenting TIPA and its extension, Multi-Token Internal Position Awareness (MTIPA), which improves models’ ability to capture character positions for accurate predictions.
@@ 2 Related Work
Tokenization methods like BPE (Sennrich, 2015; Wang et al., 2020) and WordPiece (Schuster and Nakajima, 2012) improve computational efficiency in LLMs but obscure internal character structures. Kaushal and Mahowald (2022) found that while larger models encode character-level details better, they may not explicitly understand character positions within tokens. Recent byte-level models like ByT5 (Xue et al., 2022) process raw bytes for character-level precision but require architectural changes that prevent low-cost adaptation of existing subword-based LLMs. Hybrid approaches (e.g., CANINE (Clark et al., 2022)) improve character awareness but still lack efficient positional modeling for multi-character tokens.
@@ 3.2 Multi-Token Internal Position Awareness (MTIPA)
3.3 Extended Methodology with Full-Parameter SFT
The tulu-3-sft-mixture dataset(Lambert et al., 2024) developed by AI2 enables base LLM refinement into conversational AI. We extracted vocabulary from Llama-3.1(Grattafiori et al., 2024) tokenizer, integrated all tokens into TIPA, and merged this with tulu-3-sft-mixture for full-parameter supervised fine-tuning (SFT) of Llama-3.1-8B. This produced Llama-3.1-Tulu-TIPA-8B without requiring additional character-level task datasets, maintaining the model’s inherent capabilities while enhancing character-level processing through the integrated tokenizer vocabulary.
@@ 4 Redefining the Chinese Spelling Correction Task
We redefine the Chinese Spelling Correction (CSC) task to require the model to output the positions of incorrect characters along with their corrections. For example, the sentence {CJK}UTF8gkai"业内人事东南亚正常" (translation: "Industry insiders say that withdrawing to Southeast Asia is also normal") becomes:
@@
UTF8gkai ["position": 4, "incorrect": "", "correction": ""]
Here, the character at position 4, "{CJK}UTF8gkai", is incorrect and should be corrected to "{CJK}UTF8gkai".
@@ 5 Experiments
We conducted comprehensive experiments to validate the effectiveness of TIPA and MTIPA. Our experiments are divided into three main parts: position-aware CSC, traditional CSC, and general model training and evaluation. Below, we detail the datasets, baselines, and training protocols.
@@ 5.1 Datasets
For general model training, we used tulu-3-sft-mixture, an open-source conversational AI fine-tuning dataset for the supervised fine-tuning (SFT) stage(Lambert et al., 2024). We performed full-parameter supervised fine-tuning of Llama-3.1-8B using the combined tulu-3-sft-mixture and TIPA dataset. The TIPA dataset was constructed by extracting vocabulary from the Llama-3.1 tokenizer and integrating all tokens into the TIPA format.
For general model evaluation, we used the following benchmarks: IFEVAL(Zhou et al., 2023) (instruction following evaluation with strict/loose scoring), GSM8K (Cobbe et al., 2021) (math word problem-solving), MMLU(Hendrycks et al., 2020) (massive multitask language understanding), AExams(Hardalov et al., 2020) (Arabic exam question answering), KoBEST(Jang et al., 2022) (Korean language understanding benchmark), and HumanEval(Chen et al., 2021) (Python code generation), TyDi QA(Clark et al., 2020) (Multilingual QA benchmark). Inspired by LLM The Genius Paradox(Xu and Ma, 2024), we developed multilingual datasets covering three tasks: Occurrence, Length, and Distinct.
@@ 5.2 Baseline Models
We compared our methods against several baseline models, including Pure-SFT, GPT-4o (Hurst et al., 2024), DeepSeek v2.5, ERNIE-4.0, and GLM-4-Plus, to assess the relative performance improvements. For general model evaluation, we used Llama-3.1-Tulu as a baseline.
@@ 5.3 Training Details
In Experiment 1 and Experiment 2, we fine-tuned the open-source Qwen2.5-7B (Yang et al., 2024) model using LoRA (Hu et al., 2021) to incorporate TIPA and MTIPA. LoRA allows efficient adaptation of large models without modifying the original model weights, making it suitable for our experiments.
@@
In Experiment 3, we extracted all tokens from the Llama-3.1 tokenizer vocabulary, performed TIPA operations to generate a dataset, and conducted mixed training with the tulu-3-sft-mixture dataset. Additional training details, hardware configurations, and training time can be found in Appendix A.2.
@@ 5.4.1 Results
Table 4: (Experiment 1)Results on position-based CSC using CSCD-NS test dataset, showing that TIPA-7B and MTIPA-7B outperform the baseline in all evaluated metrics. GPT-4o’s advantage stems from its character-level tokenization simplifying position identification, while other models face challenges with subword tokenization. Cross-tokenizer comparisons are less meaningful, yet previous LLMs performed poorly on this task.
@@ 5.5.1 Results
Table 5: (Experiment 2)Traditional CSC results on CSCD-NS test dataset. TIPA improves character-level detection and correction, with reverse-order TIPA showing greater gains at larger model scales.
@@ 5.5.2 Evaluation on LEMON Dataset
Table 6: (Experiment 2)Character-level F1 scores on LEMON and CSCD-NS for various domains. TIPA-7B consistently achieves higher F1, indicating better generalization.
5.6 Experiment 3: General Model Evaluation
To validate TIPA’s effectiveness beyond Chinese-specific tasks and LoRA fine-tuning, we conducted full-parameter supervised fine-tuning experiments on the Llama-3.1-8B model.
5.6.1 Results
Table 7 shows that TIPA-enhanced models maintain comparable performance on benchmarks while improving in character-sensitive tasks. The TIPA model achieves particularly strong gains in IFEVAL (strict and loose) and AExams, demonstrating enhanced instruction following capabilities.
Metric TULU TULU-TIPA
IFEVAL0-shot, strict↑ 64.88 67.84
IFEVAL0-shot, loose↑ 68.02 70.24
GSM8K8-shot↑ 74.53 74.53
MMLU5-shot↑ 65.30 65.30
AExams 38.92 40.22
KoBESTF1↑ 51.09 52.37
HumanEvalpass@1↑ 53.05 51.83
Table 7: Standard benchmark evaluation of Llama-3.1-Tulu and Llama-3.1-Tulu-TIPA-8B, showing comparable performance with TIPA-enhanced model.
Table 8 demonstrates TIPA’s effectiveness in multilingual settings, with the average F1 score improving from 47.85 to 52.81 across nine languages. The most significant gains appear in Finnish, Indonesian, and Korean, suggesting TIPA helps with non-Latin scripts.
Language F1
TULU TULU-TIPA
Arabic 67.22 67.02
Bengali 30.44 31.13
English 61.16 64.83
Finnish 49.74 60.99
Indonesian 56.76 66.64
Korean 46.16 56.37
Russian 44.07 44.95
Swahili 47.73 53.10
Telugu 27.39 30.28
Average 47.85 52.81
Table 8: TyDi QA multilingual evaluation showing improved performance with TIPA.
Table 9 confirms TIPA’s benefits for character-level tasks, with the TIPA model outperforming baseline by 3.95% on character occurrence counting, 7.11% on sentence length prediction, and achieving double the performance on distinct character counting (9.29% vs 18.74%).
Metric TULU TULU-TIPA
Occurrence 34.62 38.57
Length 24.48 31.59
Distinct 9.29 18.74
Table 9: Average character-level comparison between TULU and TIPA systems across eight languages (CN, EN, JP, KR, AR, EN-Hard, FR, RU).
@@ 6.1 Position Prediction Accuracy
Figure 3: (Experiment 1)Comparison of position prediction accuracy by character position. MTIPA-7B achieves consistently higher accuracy, especially at longer positions.
@@ 6.2 Training Dynamics
Figure 4: (Experiment 1)Training dynamics showing character-level metrics and position accuracy. MTIPA-7B achieves the highest performance ceiling.
@@ 6.3 Impact on Downstream Tasks
6.4 General Model Findings
Our general model experiments reveal several key insights:
  • TIPA integration through full-parameter SFT maintains model performance on benchmarks while improving character-level capabilities
  • The multilingual evaluation shows TIPA’s benefits extend beyond Chinese to other scripts and languages
  • Character-level tasks demonstrate significant improvements, particularly in counting distinct characters
  • The approach scales effectively to larger models (8B parameters) without performance degradation
These results suggest that TIPA’s benefits are not limited to Chinese-specific tasks or LoRA fine-tuning, but represent a general improvement in LLMs’ character-level understanding.
@@ 7 Conclusion
We introduced Token Internal Position Awareness (TIPA) and Multi-Token Internal Position Awareness (MTIPA) to enhance large language models’ (LLMs) ability to accurately predict character positions within tokens. Our experiments demonstrate TIPA’s effectiveness across three scenarios: position-aware Chinese spelling correction, traditional CSC tasks, and both training and evaluation of general models. The method shows particular strength in:
1. Improving position prediction accuracy in Chinese text (up to 8.07% absolute gain) 2. Enhancing traditional CSC performance (up to 5.64% F1 improvement) 3. Boosting multilingual and character-level capabilities without sacrificing benchmark performance 4. Scaling effectively to larger models through both LoRA and full-parameter fine-tuning
By training LLMs on reverse character prediction tasks using the tokenizer’s vocabulary, TIPA, and MTIPA effectively address the limitations imposed by tokenization methods like Byte-Pair Encoding (BPE) while maintaining compatibility with existing systems.
@@ 8 Limitations
  • Directly mixing TIPA datasets during general model SFT training may induce a bias toward shorter text sequence generation when using excessive data proportions. While this can be mitigated by setting minimum token constraints, more fundamental solutions through pretraining or reinforcement learning stages warrant further exploration.
9 Acknowledgments
This work was supported by the "Major Special Project for Technological Innovation and Application Development" (Grant No. CSTB2024TIAD-STX0036) from the Chongqing Municipal Science and Technology Bureau. We sincerely appreciate the insightful comments and constructive suggestions from the reviewers, which significantly improved the quality and rigor of this manuscript. We are especially grateful to Dr. Yixian Shen, a postdoctoral researcher at the University of Amsterdam, for generously providing computational resources and invaluable guidance during the large-scale model training phase. Their expertise and support were instrumental in advancing this research.
@@ A.2 Implementation Details
A.2.1 TIPA Training Methodology
We implemented two approaches for TIPA training:
  • Unpruned TIPA: We first extracted all vocabulary from the tokenizer and filtered out tokens that could not be properly parsed as UTF-8. We then applied the TIPA method to create the TIPA dataset, which was mixed with the CSC dataset and shuffled before LoRA fine-tuning.
  • Pruned TIPA: We tokenized a large CSC dataset and intersected these tokens with the tokenizer’s vocabulary. This pruned approach resulted in a TIPA dataset containing primarily Chinese characters and a small portion of other language tokens, while still covering nearly all Chinese characters. This dataset was then mixed with the CSC dataset and shuffled before LoRA fine-tuning.
A.2.2 MTIPA Training Methodology
For MTIPA implementation, we:
  • Selected either the unpruned or pruned TIPA method to create the base TIPA dataset
  • Randomly sampled incorrect source texts from the dataset
  • Decomposed the sampled strings using the same reverse decomposition task as for individual tokens
  • Mixed the resulting MTIPA dataset with the CSC and TIPA datasets
  • Shuffled the combined dataset before LoRA fine-tuning
This combined training approach ensures that the model learns both token-internal structures and their application in context while maintaining the efficiency of standard fine-tuning procedures.
@@ A.2.2 MTIPA Training Methodology
A.2.3 Full-Parameter SFT Configuration
For general model training without LoRA adaptation, we implemented complete parameter fine-tuning using the configuration shown in Table 15:
Configs Values
Devices 4 NVIDIA H100 GPUs (80GB)
Batch size 32 (gradient accumulation)
Learning rate $5\times 10^{-6}$
Epochs 2
Table 15: Full-parameter SFT configurations. All unspecified parameters align with TULU training settings. Training completed in 3d 8h 10m 26s.
@@ A.3 Datasets
We utilized several datasets for training and evaluation. The datasets and their characteristics are summarized in Tables 16 and 17. The total number of unique tokens after deduplication across all datasets is 24,994.
@@ A.7 Loss Analysis
Figure 6: (Experiment 2)Training and validation loss comparison across different methods in the traditional CSC task. The TIPA-7B model exhibits the fastest reduction in loss during training, indicating more efficient learning and better generalization capabilities compared to the baseline.